# Exploring selection biases in Fast Radio Burst (FRB) dispersion-galaxy cross-correlations with magnetohydrodynamical simulations

### v1

@@ Abstract
The dispersion of fast radio bursts (FRBs) in conjunction with their redshifts can be used as powerful probes of the distribution of extragalactic plasma, and with a large enough sample, the free-electron–galaxy power spectrum $P_{eg}$ can be measured by cross-correlating FRB dispersions with galaxy positions. However, a precise measurement of $P_{eg}$ requires a careful investigation of the selection effects—the fact that the probability of both observing the FRB dispersion measure and obtaining a host galaxy redshift depends on observed properties of the FRB and its host. We use ray tracing simulations with IllustrisTNG300-1 to investigate the impact of expected observational selection effects on FRB dispersion–galaxy angular cross-correlations with a sample of 3000 FRBs at redshift range of $0.3\leq z\leq 0.4$ Our results show that cross-correlations with such an FRB sample are robust to properties of the FRB host galaxy: this includes DM contributions from the FRB host and optical followup selection effects biased against FRBs with dim galaxy hosts. We also find that such cross-correlations are robust to DM dependent and scattering selection effects specific to the CHIME/FRB survey. However, a DM dependent selection effect that cuts off the 10% most dispersed FRB at a fixed redshift shell can bias the amplitude of the cross-correlation signal by over 50% at angular scales of $\sim 0.1^{\circ}$ (corresponding to Mpc physical scales). Our findings highlight the importance of both measuring and accounting for selection effects present in existing FRB surveys as well as mitigating DM dependent selection effects in the design of upcoming FRB surveys aiming to use FRBs as probes for large-scale structure.
1 Introduction
@@
Fast Radio Bursts (FRBs) are extragalactic millisecond-duration radio transients originating at cosmological redshifts. They are characterized by their large dispersive delays, parameterized via dispersion measures (DM), where the dispersion measure (DM) is directly proportional to the integrated free electron density along the line of sight. Since the majority of the Universe’s baryonic matter is in the form of a diffuse plasma outside of galaxies, FRB dispersion is an excellent tracer of baryons. Unlike other tracers of baryonic matter, such as X-ray observations and the thermal Sunyaev-Zeldovich effect, which have a temperature-dependent signal, FRB DMs probe hot and cold ionized gas equally. As such, FRBs have emerged as promising cosmological probes that can help solve a range of astrophysical problems that are sensitive to the distribution of baryonic matter (Lorimer et al., 2007; Thornton et al., 2013; McQuinn, 2014; Macquart et al., 2020), from improving galaxy-cluster kinetic SZ measurements (Madhavacheril et al., 2019) to constraining models of astrophysical feedback (Medlock et al., 2024; Sharma et al., 2025).
@@
While to date thousands of FRBs have been detected (CHIME/FRB Collaboration et al., 2021a), only $\sim 100$ have been localized with the angular resolution necessary to identify their host galaxies and obtain redshifts (Hallinan et al., 2019; Bhandari et al., 2020; FRB Collaboration et al., 2025a). Nevertheless, enough FRB redshifts have been measured to enable studies of the extragalactic electron density distribution through the FRB line of sight via “one-point” statistics of DM distributions. This typically requires partitioning the total dispersion measure of an FRB into its constituent parts (e.g. a contribution from the FRB host galaxy, a cosmic contribution from the CGMs intervening halos and the IGM, and a contribution from the Milky Way) (Connor et al., 2024; Sharma et al., 2025) that assume a parameterization for each component based on hydrodynamical simulations. Parameters of interest are then used to probe a given astrophysical process; for instance, the so called “F-parameter” originally used by Macquart et al. (2020) and then again by Connor et al. (2024) to constrain the fraction of all baryons in collapsed halos, parametrizes the variance of the cosmic contribution of the DM as $\sigma_{DM,\rm{cosmic}}=Fz^{-1/2}$, while Sharma et al. (2025) proposes parameterizing the variance as a redshift dependent log-normal distribution which can in turn be used to constrain the strength of different astrophysical feedback processes.
@@
In general there are two main disadvantages to using one-point statistics of the DM. The first is that the precision of the DM-budgeting procedure is limited by the accuracy of the (likely overly simplistic) underlying parametrization of each DM component, which can be motivated but not verified by simulations. The second is sensitivity to selection effects—the fact that the probability of detecting an FRB depends on its observed properties—which can bias the baryon distribution as measured from FRB DMs. For instance, both the ASKAP and CHIME/FRB search pipelines are insensitive to over 50% of the expected FRB population at DM$>$1000pc/cm-3 (Shannon et al., 2018; Shin et al., 2023; Merryfield et al., 2023). If, as a simple example, this preferentially removes FRB detections that intersect halos with a dense electron density, a measurement of the distribution of baryons in collapsed halos using the one-point statistic would be an underestimate.
@@
In this work, we demonstrate through our ray tracing simulations which selection biases cross-correlations are the most sensitive to, along with the rough magnitudes of the effects. In Section 2 we describe the details of our cross-power spectrum estimation and ray tracing simulations, and in Section 3 we present and discuss the results of our different simulated selection biases. We conclude in Section 4.
2 Preliminaries and Methods
2.1 The DM-galaxy cross-correlation power spectrum
@@
[…]
where $n_{e}$ denotes the comoving electron number density and $z$ is the redshift. Therefore, the average electron number density $\bar{n}_{e}(z)$ gives us $\overline{\mathrm{DM}}(z)$, the expected DM of an FRB at redshift $z$; this is the well-known Macquart relation (Macquart et al., 2020). Given a catalog of FRBs, we can then define the DM overdensity field as
@@
[…]
$C^{Dg}_{\ell}=n_{e,0}\frac{1+z_{g}}{\chi_{g}^{2}}P_{eg}(k=\ell/\chi_{g},z_{g}).$ (5)
where $P_{eg}$ in turn probes the distribution of baryonic matter around galaxies.
2.2 Power Spectrum Estimation
@@
In a realistic survey, the full covariances $\mathbf{D}$ and $\mathbf{G}$ can be difficult to estimate. Even if they can be obtained, the computational cost for Equation 13 is often too high. Works such as Alonso et al. (2019) have shown that simplifications by assuming the covariances $\mathbf{D}^{-1}$ and $\mathbf{G}^{-1}$ are diagonal allow Equations 12 and 13 to be computed efficiently while keeping the estimator close to optimal in most cases. One simple way to estimate $\mathbf{D}^{-1}$ and $\mathbf{G}^{-1}$ is to have some estimated variance $\sigma^{-2}$ along the diagonal for pixels with at least one FRB, and 00 for pixels with no FRBs, corresponding to a limit of infinite noise (i.e. the DM is not measured). However, any constant that multiplies $\mathbf{D}^{-1}$ or $\mathbf{G}^{-1}$ will cancel out between the inverse Fisher matrix and the numerator in Equation 12. Therefore, we can simply take the diagonal of $\mathbf{D}^{-1}$ and $\mathbf{G}^{-1}$ to be the survey mask:
@@
where $C^{DD}_{\lambda}$ and $C^{gg}_{\lambda}$ are the auto power spectrum for the DM and galaxy overdensity field, respectively. However, as we discuss further in Section 2.4, the Gaussianity assumption on which this equation relies is unlikely to be valid on small scales due to non-linear structure growth. We therefore estimate the errors of our cross-power spectrum empirically, by computing it over multiple realizations in many independent sky patches (see Section 2.3).
2.3 Ray Tracing with IllustrisTNG
@@
[…]
@@
The larger the distance to a given FRB, the more times the simulation box must be stacked and therefore the more likely it is that a given line of sight will intersect with the same structure more than once. Equivalently, the larger the FRB redshift, the smaller the allowed sky regions become, therefore requiring a compromise between these quantities. We leave the details of these computations to Appendix B. The result is summarized in Figure 1, which shows the allowed sky regions for our choice of $z_{\max}=0.4$ and pixel resolution $0.0008\,\mathrm{rad}$, along with the $48$ $0.1\,\mathrm{rad}\times 0.1\,\mathrm{rad}$ square sky patches within the allowed regions. There are $15625125^{2}=$ 15625 pixels per sky region.
Figure 1: The $48$ square sky regions (shaded gray) used in this study (out to $z_{\max}=0.4$) for independent realizations of the DM and galaxy field. The regions are chosen from the “good” regions of the sky (white) that avoid spurious geometrical effects from the simulation’s periodic boundary conditions.
@@
Finally, the DM-galaxy cross-correlation requires a catalog of foreground galaxies. Because there are far fewer galaxies than simulation particles, it is not necessary to bin the galaxy map as we did for electrons. IllustrisTNG provides galaxies in the form of a “group catalog” at each snapshots, containing a list of halos and subhalos. Galaxies are taken to be subhalos flagged as galaxies (“SubhaloFlag=0”) with nontrivial stellar content ($M_{g}<0$). We build the light cone of galaxies in a similar fashion as our ray tracing procedure, stacking simulation boxes and evolving the cone in tomographic slices corresponding to each snapshot. We identify the redshift range each snapshot, computed as the halfway point (in comoving distance) to neighboring snapshots. Then, for each snapshot $i$, we create a galaxy catalog corresponding to that slice by taking all galaxies with $\chi(z_{\min,i})\leq|\vec{x}_{g}-\vec{x}_{\text{obs}}|<\chi(z_{\max,i})$, where $\vec{x}_{g}$ and $\vec{x}_{\text{obs}}$ are the positions of the galaxy and observer, respectively. We save the galaxies’ center-of-mass coordinates, along with other properties retrieved from IllustrisTNG (stellar mass, star formation rate (SFR), and $g$-band absolute magnitude), from which we straightforwardly compute the projected sky position and apparent magnitude. The end product is a catalog of galaxies within the cone of each sky region up to a maximum redshift.
@@
2.4 Fiducial cross-power spectrum
@@
In all our experiments, we simulate $N=3000$ FRBs within a redshift slice $0.3<z_{\text{host}}\leq 0.4$ for each sky region shown in Figure 1. We draw our FRBs randomly from galaxies weighted by the galaxy’s SFR under the simple model that FRBs trace the star formation rate:
@@
We also note that the error bars (standard deviation over all $C^{Dg}_{\ell}$) placed on our power spectra are the error bars corresponding to a survey of 3000 FRBs, not the uncertainties on the reported mean simulated power spectrum measurement (the standard error) which are a factor of $\sqrt{48}$ smaller. The factor of $\sqrt{48}$ can be understood as follows: the total contribution to the cross correlation noise in an ideal survey of FRBs with redshifts (i.e. the Maquart relation is removed) can be decomposed into (a) sample/cosmic variance (b) FRB host galaxy/local environment noise and (c) foreground galaxy shot noise. In the case of our simulation where all foreground galaxies are used in the cross correlation, the galaxy shot noise can be ignored, leaving contributions (a) and (b). While each of the 5 trials randomizing the choice of background FRBs reduce the contribution of (b), the large-scale structure sample variance is the same for each trial within a given sky region, meaning we have only 48 effective realizations of (a).
@@
3 Impact of Selection Effects
3.1 Insensitivity to FRB and host galaxy properties

### v2

# Exploring selection biases in FRB dispersion–galaxy cross-correlations with magnetohydrodynamical simulations

@@ Abstract
The dispersion measure (DM) of Fast Radio Bursts can be used as powerful probes of the distribution of extragalactic plasma. With a large enough sample, the free-electron–galaxy power spectrum $P_{eg}$ can be measured by cross-correlating FRB DMs with galaxy positions. However, a precise measurement of $P_{eg}$ requires a careful investigation of selection effects: the probability of both observing the FRB DM and obtaining a host galaxy redshift depends on their properties. We ray trace through the magnetohydrodynamic simulation IllustrisTNG to investigate the impact of expected observational selection effects on FRB dispersion–galaxy angular cross-correlations with a sample of 3000 FRBs at $0.3\leq z\leq 0.4$ . Our results show that cross-correlations with such an FRB sample are robust to properties of the FRB host galaxy: this includes DM contributions from the FRB host and optical follow-up selection effects. We also find that such cross-correlations are robust to DM-dependent and scattering selection effects specific to the CHIME/FRB survey. However, a DM-dependent selection effect that cuts off the 10% most dispersed FRB at a fixed redshift shell can bias the amplitude of the cross-correlation signal by over 50% at angular scales of $\sim 0.1^{\circ}$, corresponding to $\sim$ Mpc physical scales. Our findings highlight the importance of both measuring and accounting for selection effects present in existing FRB surveys, as well as mitigating DM-dependent selection effects in the design of upcoming FRB surveys aiming to probe large-scale structure with FRBs.
I Introduction
@@
Fast Radio Bursts (FRBs) are extragalactic millisecond-duration radio transients originating at cosmological redshifts. They are characterized by their large dispersive delays, parameterized via DMs, where the DM is directly proportional to the integrated free electron density along the line of sight. Since the majority of the Universe’s baryonic matter is in the form of a diffuse plasma outside of galaxies, FRB dispersion is an excellent tracer of baryons (Lorimer et al., 2007; Petroff et al., 2022). Unlike other tracers of baryonic matter, such as X-ray observations (e.g. Pen, 1999; Nicastro et al., 2018) and the thermal Sunyaev-Zeldovich effect (e.g. Tanimura et al., 2019; de Graaff et al., 2019), which have a signal that is strongly temperature-dependent, FRB DMs are sensitive to every free electron along the line of sight. As such, FRBs have emerged as promising cosmological probes that can help solve a range of astrophysical problems that are sensitive to the distribution of baryonic matter (Lorimer et al., 2007; Thornton et al., 2013; McQuinn, 2014; Macquart et al., 2020), from improving galaxy-cluster kinetic SZ measurements (Madhavacheril et al., 2019) to constraining models of astrophysical feedback (Medlock et al., 2024; Sharma et al., 2025).
@@
While to date thousands of FRBs have been detected (CHIME/FRB Collaboration et al., 2021b), only $\sim 100$ have been localized with the angular resolution necessary to identify their host galaxies and obtain redshifts (Hallinan et al., 2019; Heintz et al., 2020; Bhandari et al., 2020; Gordon et al., 2023; Law et al., 2024; Bhardwaj et al., 2024; Sharma et al., 2024; FRB Collaboration et al., 2025a). Nevertheless, enough FRB redshifts have been measured to enable studies of the extragalactic electron density distribution through the FRB line of sight via “one-point” statistics of DM distributions. This typically requires partitioning the total dispersion measure of an FRB into its constituent parts (e.g. a contribution from the FRB host galaxy, a cosmic contribution from intervening halos and the intergalactic medium, and a contribution from the Milky Way) (Connor et al., 2024; Sharma et al., 2025) that assume a parameterization for each component informed by observations or hydrodynamical simulations. Parameters of interest can then used to probe a given astrophysical process. For instance, the so called “F-parameter”, originally used by Macquart et al. (2020) to constrain the fraction of all baryons in collapsed halos, parametrizes the variance of the cosmic contribution of the DM at a given redshift as $\sigma_{\text{DM, cosmic}}(z)=Fz^{-1/2}$. Connor et al. (2024) and Sharma et al. (2025) use simulation-based inference to constrain the strength of different astrophysical feedback processes by modelling the DM distribution at a given $z$ as a log-normal.
@@
In general there are two main disadvantages to using one-point statistics of the DM. The first is that the precision of the DM-budgeting procedure is limited by the accuracy of the (likely overly simplistic) underlying parametrization of each DM component, which can be motivated but not verified by simulations. The second is sensitivity to selection effects—the fact that the probability of detecting an FRB depends on its observed properties (James et al., 2022a, b)—which can bias the baryon distribution as measured from FRB DMs. For instance, both the ASKAP and CHIME/FRB search pipelines are insensitive to over 50% of the expected FRB population at DM >> 1000 pc/cm-3 (Shannon et al., 2018; Shin et al., 2023; Merryfield et al., 2023). If, as a simple example, this preferentially removes FRB detections that intersect halos with a dense electron density, a measurement of the distribution of baryons in collapsed halos using the one-point statistic would be an underestimate.
@@
However, with upcoming next generation radio telescopes such as CHORD, DSA 2000, and BURSTT, $\sim 10^{4}-10^{6}$ FRBs are expected to be localized to their host galaxies in the next decade (Vanderlinde et al., 2019; Hallinan et al., 2019; Lin et al., 2022), allowing for studies on large-scale structure with FRB redshifts to be extended from the existing one-point measurements to two-point cross-correlations.
@@
In this work, we demonstrate through our ray tracing simulations which selection biases cross-correlations are the most sensitive to, along with the rough magnitudes of the effects. In Section II we describe the details of our cross-power spectrum estimation and ray tracing simulations, and in Section III we present and discuss the results of our different simulated selection biases. We conclude in Section IV.
II Preliminaries and Methods
II.1 The DM-galaxy cross-correlation power spectrum
@@
$\mathrm{DM}(\chi,{\vec{\theta}})=\int_{0}^{\chi}\,d\chi^{\prime}\,n_{e}(\chi^{\prime},{\vec{\theta}})(1+z(\chi^{\prime})),$ (1)
where $n_{e}$ denotes the comoving electron number density and $z$ is the redshift. Therefore, the average electron number density $\bar{n}_{e}(z)$ gives us $\overline{\mathrm{DM}}(z)$, the expected DM of an FRB at redshift $z$; this is the well-known Macquart relation (Macquart et al., 2020). Given a catalog of FRBs, we can then define the DM perturbation as
@@
where $\langle\ldots\rangle$ denotes an ensemble average, and $\tilde{d}_{lm}(z_{host})$ and $\tilde{g}_{lm}(z_{g})$ correspond to the spherical harmonic transform (SHT) of the DM and galaxy number overdensity fields at their respective redshifts (we will suppress the redshift dependence for simplicity in our following discussion). The cross-correlation $C^{Dg}_{\ell}$ between the DM field of background FRBs and a foreground galaxy field at redshift $z_{g}$ is significant because it provides a direct measuremment of the galaxy-electron 3D power spectrum $P_{eg}$ via the relation (Madhavacheril et al., 2019)
$C^{Dg}_{\ell}=n_{e,0}\frac{1+z_{g}}{\chi_{g}^{2}}P_{eg}\left(k=\frac{\ell+1/2}{\chi_{g}},z_{g}\right),$ (5)
$P_{eg}$, in turn, probes the distribution of baryonic matter around galaxies. Note that Equation 5 is derived using the Limber approximation, which is valid up to very large scales (small $\ell$), at which we are simultaneously limited by the finite simulation periodic box size.
II.2 Power Spectrum Estimation
@@
In a realistic survey, the full covariances $\mathbf{D}$ and $\mathbf{G}$ can be difficult to estimate. Even if they can be obtained, the computational cost for Equation 13 is often too high. Works such as Alonso et al. (2019) have shown that simplifications by assuming the covariances $\mathbf{D}^{-1}$ and $\mathbf{G}^{-1}$ are diagonal allow Equations 12 and 13 to be computed efficiently while keeping the estimator close to optimal in most cases. One simple way to estimate $\mathbf{D}^{-1}$ and $\mathbf{G}^{-1}$ is to have some estimated variance $\sigma^{-2}$ along the diagonal for pixels with at least one FRB, and 0 for pixels with no FRBs, corresponding to a limit of infinite noise (i.e. the DM is not measured). However, any constant that multiplies $\mathbf{D}^{-1}$ or $\mathbf{G}^{-1}$ will cancel out between the inverse Fisher matrix and the numerator in Equation 12. Therefore, we can simply take the diagonal of $\mathbf{D}^{-1}$ and $\mathbf{G}^{-1}$ to be the survey mask:
@@
where $C^{DD}_{\lambda}$ and $C^{gg}_{\lambda}$ are the auto power spectrum for the DM and galaxy overdensity field, respectively. However, as we discuss further in Section II.4, the Gaussianity assumption on which this equation relies is unlikely to be valid on small scales due to non-linear structure growth. We therefore estimate the errors of our cross-power spectrum empirically, by computing it over multiple realizations in many independent sky patches (see Section II.3).
II.3 Ray Tracing with IllustrisTNG
@@
Due to the memory limitations of the JupyterLab workspace, which limits the size (total number of pixels) of the output gridded map that can be held in memory, we choose a bin resolution of $500\,\mathrm{ckpc}/h$. We note that our pixel resolution of 500 ckpc/$h$ places an upper limit of $k\lesssim\sim 6\,h\,\mathrm{Mpc}^{-1}$ on the scales at which we can resolve $P_{eg}$, whereas the finite periodic box size sets a lower limit of $k\gtrsim 0.02$. For our fiducial $0.2<z_{g}<0.3$, this corresponds to $20\lesssim\ell\lesssim 5000$.
@@
The larger the distance to a given FRB, the more times the simulation box must be stacked and therefore the more likely it is that a given line of sight will intersect with the same structure more than once. Equivalently, the larger the FRB redshift, the smaller the allowed sky regions become, therefore requiring a compromise between these quantities. We leave the details of these computations to Appendix B. The result is summarized in Figure 1, which shows the allowed sky regions for our choice of $z_{\max}=0.4$ and pixel resolution $0.0008\,\mathrm{rad}$, along with the 4848 $0.1\,\mathrm{rad}\times 0.1\,\mathrm{rad}$ square sky patches within the allowed regions. There are $125^{2}=15625$ pixels per sky region.
Figure 1: The 4848 square sky regions (shaded gray) used in this study (out to $z_{\max}=0.4$) for independent realizations of the DM and galaxy field. The regions are chosen from the “good” regions of the sky (white) that avoid spurious geometrical effects from the simulation’s periodic boundary conditions.
@@
Finally, the DM-galaxy cross-correlation requires a catalog of foreground galaxies. Because there are far fewer galaxies than simulation particles, it is not necessary to bin the galaxy map as we did for electrons. IllustrisTNG provides galaxies in the form of a “group catalog” at each snapshots, containing a list of halos and subhalos. Galaxies are taken to be subhalos flagged as galaxies (“SubhaloFlag=0”) with nontrivial stellar content ($M_{g}<0$). Note that while this results in a halo mass cutoff that is lower than realistic, imposing realistic apparent magnitude cuts to our mock galaxy catalog was not found to significantly change our results, with only a modest increase in the error bars.
With the same simulation volume as the DM map, we build the light cone of galaxies in a similar fashion as our ray tracing procedure, stacking simulation boxes and evolving the cone in tomographic slices corresponding to each snapshot. We identify the redshift range each snapshot, computed as the halfway point (in comoving distance) to neighboring snapshots. Then, for each snapshot $i$, we create a galaxy catalog corresponding to that slice by taking all galaxies with $\chi(z_{\min,i})\leq|\vec{x}_{g}-\vec{x}_{\text{obs}}|<\chi(z_{\max,i})$, where $\vec{x}_{g}$ and $\vec{x}_{\text{obs}}$ are the positions of the galaxy and observer, respectively. We save the galaxies’ center-of-mass coordinates, along with other properties retrieved from IllustrisTNG (stellar mass, star formation rate (SFR), and $g$-band absolute magnitude), from which we straightforwardly compute the projected sky position and apparent magnitude. The end product is a catalog of galaxies within the cone of each sky region up to a maximum redshift.
@@
II.4 Fiducial cross-power spectrum
@@
In all our experiments, we simulate $N=3000$ FRBs within a redshift slice $0.3<z_{\text{host}}\leq{z_{\mathrm{\max}}}=0.4$ for each sky region shown in Figure 1. We draw our FRBs randomly from galaxies weighted by the galaxy’s SFR under the simple model that FRBs trace the star formation rate:
@@
We also note that the error bars (standard deviation over all $C^{Dg}_{\ell}$) placed on our power spectra are the error bars corresponding to a survey of 3000 FRBs, not the uncertainties on the reported mean simulated power spectrum measurement (the standard error) which are a factor of $\sqrt{48}$ smaller. The factor of $\sqrt{48}$ can be understood as follows: the total contribution to the cross correlation noise in an ideal survey of FRBs with redshifts (i.e. the Maquart relation is removed) can be decomposed into (a) sample/cosmic variance (b) FRB host galaxy/local environment noise and (c) foreground galaxy shot noise. As discussed in Section II.3, we use a large sample of all foreground galaxies, and thus the galaxy shot noise is negligible, leaving contributions (a) and (b). While each of the 5 trials randomizing the choice of background FRBs reduce the contribution of (b), the large-scale structure sample variance is the same for each trial within a given sky region, meaning we have only 48 effective realizations of (a).
@@
III Impact of Selection Effects
III.1 Insensitivity to FRB and host galaxy properties